# Nearly Optimal Sample Complexity of Offline KL-Regularized Contextual Bandits under Single-Policy Concentrability

### v1

@@ 2.2 Review of Existing Results
In this subsection, we discuss the direct adaptation of previous results to our setting and demonstrate that their results cannot imply an $\widetilde{O}(\epsilon^{-1})$ sample complexity without all-policy concentrability. Specifically, Xiong et al. (2024) obtained a performance gap upper bound under linear function approximation
@@
2.3 Algorithm
In this subsection, we present an offline bandit algorithm, KL-PCB, for KL-regularized contextual bandits in Algorithm 1. KL-PCB first leverages least-square estimator to find a function ${\bar{f}}\in\mathcal{F}$ that minimizes its risk on the offline dataset. In Zhao et al. (2024), such ${\bar{f}}$ is directly applied to construct the estimated policy. In contrast, we construct a pessimistic estimator of $f^{*}$ following the well-known pessimism principle in offline RL (Jin et al., 2021).
Specifically, we consider the bonus term defined as
$\Gamma_{n}(s,a)=\beta D_{\mathcal{F}}\big{(}(s,a),\pi^{\mathsf{ref}}\big{)},\forall(s,a)\in{\mathcal{S}}\times\mathcal{A};$ (2.3)
where
$\beta=\sqrt{128\log\big{(}2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta\big{)}/3n+18\epsilon_{c}}.$ (2.4)
We then obtain our pessimistic estimation ${\widehat{f}}$ by setting ${\widehat{f}}={\bar{f}}-\Gamma_{n}$, which is less than $f^{*}$ with high probability. After obtaining the pessimistic estimation, KL-PCB output the policy $\widehat{\pi}$, which maximizes the estimated objective
$\widehat{J}(\pi)=\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg{[}{\widehat{f}}(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]},$
the maximizer of which is the counterpart of 2.2, i.e.,
$\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{f}}(s,a)\big{)}.$
Algorithm 1 Offline KL-Regularized Pessimistic Contextual Bandit (KL-PCB)
0:  regularization $\eta$, reference policy $\pi^{\mathsf{ref}}$, function class $\mathcal{F}$, offline dataset $\mathcal{D}=\{(s_{i},a_{i},r_{i})\}_{i=1}^{n}$
1:  Compute the least square estimation of reward function
${\bar{f}}\in\mathop{\mathrm{argmin}}_{f\in\mathcal{F}}\sum_{(s_{i},a_{i},r_{i})\in\mathcal{D}}\big{(}f(s_{i},a_{i})-r_{i}\big{)}^{2}$
2:  Let ${\widehat{f}}\leftarrow{\bar{f}}-\Gamma_{n}$, where $\Gamma_{n}$ is the bonus term in 2.3
2:  $\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{f}}(s,a)\big{)}$
2.4 Theoretical Results
Theorem 2.8.
Under Assumption 2.7, for sufficiently small $\epsilon\in(0,1)$, if we set $\Gamma_{n}$ as in (2.3), then $n=\widetilde{O}(\eta D_{\pi^{*}}^{2}\epsilon^{-1})$ suffices to guarantee the output policy $\widehat{\pi}$ of Algorithm 1 to be $\epsilon$-optimal with probability at least $1-\delta$.
Remark 2.9.
Previously, Zhao et al. (2024) achieved an $\widetilde{O}(\epsilon^{-1})$ sample complexity under Assumption 2.6. As a comparison, KL-PCB achieves the same $\widetilde{O}(\epsilon^{-1})$ sample complexity but only requiring Assumption 2.7, which is weaker than Assumption 2.6
We also provide the sample complexity lower bound of KL-regularized contextual bandits in the following theorem.
Theorem 2.10.
For any $\epsilon\in(0,1),\eta>0$, and algorithm Alg, there is a KL-regularized contextual bandit instance with $O(1)$ single-policy concentrability such that Alg requires at least $\Omega\big{(}\min\{{\eta\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon},{\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon^{2}}\}\big{)}$ samples to return an $\epsilon$-optimal policy.
Remark 2.11.
Theorem 2.10 is concluded by plugging $D^{2}\geq D^{2}_{\pi^{*}}$ to Theorem 3.1 of Zhao et al. (2024) and thus the sample complexity lower bound also holds for the KL-regularized contextual bandit instance with $O(1)$ all-policy concentrability.
Remark 2.12.
Theorem 2.10 shows that when $\epsilon$ is sufficiently small, any algorithm for offline KL-regularized contextual bandits requires at least $\widetilde{\Omega}(\epsilon^{-1})$ samples to output an $\epsilon$-optimal policy. This sample complexity lower bound matches the sample complexity upper bound in Theorem 2.8, indicating that KL-PCB is nearly optimal.
2.5 Proof of Theorem 2.8
In this section, we finish the proof of Theorem 2.8. At a high level, if we consider the regularized objective 1.1 multi-arm bandits, then $P\mapsto{\mathsf{KL}}\left(P\|Q\right)$ is $1$-strongly convex w.r.t. ${\mathsf{TV}}\left(\cdot\|\cdot\right)$ (Polyanskiy and Wu, 2024, Exercise I.37), and thus $J(\pi)$ is strongly concave. Therefore, $J(\pi^{*})-J(\widehat{\pi})$ is possible to be of the order $[{\mathsf{TV}}\left(\pi^{*}\|\widehat{\pi}\right)]^{2}\approx\widetilde{O}(n^{-1})$, pretending that $\pi^{*}$ is the unconstrained maximizer. This intuition guides our analysis for contextual bandits.
$\mathcal{E}(\delta)\coloneqq\bigg{\{}\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\Big{[}\big{|}{\bar{f}}-f^{*}\big{|}-\Gamma_{n}\Big{]}(s,a)\leq 0\bigg{\}},$ (2.5)
where $\Gamma_{n}$ is defined in (2.3). Event $\mathcal{E}(\delta)$ holds indicates that the least square estimation ${\bar{f}}$ obtained in Line 1 of Algorithm 1 does not deviate too much from the true function $f^{*}$. More specifically, we have the following lemma, whose proof are deferred to Please refer to Appendix A.1.
Lemma 2.13.
The following covariance-type observation is the first pivot.
Lemma 2.14.
Remark 2.15.
While this lemma is elementary, to the best of our knowledge, we are the first to isolate this structure from our non-standard analysis of offline RL, from which the sharp upper bound is derived.

### v2

@@ A.1 Previous Attempts on Understanding KL-regularized RL
Here, we briefly discuss the direct adaptation of previous sample complexity results (with respect to KL-regularized objectives) to our setting and demonstrate the reason why theirs cannot imply an $\widetilde{O}(\epsilon^{-1})$ sample complexity without all-policy concentrability. Specifically, Xiong et al. (2024) obtained a performance gap upper bound under linear function approximation
@@
Appendix B Missing Proofs from Section 2
B.1 Proof of Lemma 2.12
We first provide the following lemmas of concentration.
Lemma B.1 (Zhao et al. 2024, Lemma C.1).
For any policy $\pi$ and state-action pairs $\{(s_{i},a_{i})\}_{i=1}^{m}$ generated i.i.d. from $\rho\times\pi$, and $\epsilon_{c}<1$, with probability at least $1-\delta$, for any $g_{1}$ and $g_{2}$ we have
$\mathbb{E}_{\rho\times\pi}\big{[}\big{(}g_{1}(s,a)-g_{2}(s,a)\big{)}^{2}\big{]}\leq\frac{2}{n}\sum_{i=1}^{n}\big{(}g_{1}(s_{i},a_{i})-g_{2}(s_{i},a_{i})\big{)}^{2}+\frac{32}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+10\epsilon_{c},$
where $\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})$ is the $\epsilon_{c}$-covering number of ${\mathcal{G}}$.
Lemma B.2 (Zhao et al. 2024, Lemma C.2).
For arbitrary policy $\pi$ and dataset $\{(s_{i},a_{i},r_{i})\}_{i=1}^{m}$ generated i.i.d., from the product of $\pi$, $\rho$ and the Bradley-Terry Model; let ${\bar{g}}$ be the least square estimator of ${g^{*}}$, then for any $10<\epsilon_{c}<$ 1 and $\delta>0$, with probability at least $1-\delta$ we have
$\sum_{i=1}^{n}\big{(}{\bar{g}}(s_{i},a_{i})-{g^{*}}(s_{i},a_{i})\big{)}^{2}\leq 16\log(a\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+4n\epsilon_{c}.$
Proof of Lemma 2.12.
We have the following inequality
$\big{(}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{)}^{2}$ $=\frac{\big{(}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{)}^{2}}{\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{)}^{2}\big{]}}\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{)}^{2}\big{]}$
$\leq\sup_{g_{1},g_{2}\in{\mathcal{G}}}\frac{\big{(}g_{1}(s,a)-g_{2}(s,a)\big{)}^{2}}{\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}g_{1}(s,a)-g_{2}(s,a)\big{)}^{2}\big{]}}\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{)}^{2}\big{]}$
$=D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{)}^{2}\big{]},$ (B.1)
where the inequality holds by taking supremum to $g_{1},g_{2}\in{\mathcal{G}}$. Now we have
$\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{)}^{2}\big{]}$ $\leq\frac{2}{n}\sum_{i=1}^{n}\big{(}{\bar{g}}(s_{i},a_{i})-{g^{*}}(s_{i},a_{i})\big{)}^{2}+\frac{32}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+10\epsilon_{c}$
$\leq\frac{2}{n}\big{[}16\log(\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+4n\epsilon_{c}\big{]}+\frac{32}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+10\epsilon_{c}$
$=\frac{128}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+18\epsilon_{c},$ (B.2)
where the first inequality holds due to Lemma B.1 and second holds due to Lemma B.2. Plugging (B.1) into (B.1) and setting $\epsilon_{c}=O(n^{-1})$ complete the proof. ∎
B.2 Proof of Lemma 2.13